Circular No. 5881

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

CH CYGNI J. P. Aufdenberg, K. D. Gordon, and B. W. Bopp, Ritter Observatory,

University of Toledo, write: “The symbiotic star CH Cyg has shown an abrupt change in the emission profiles of Hα, He I 587.5-nm, and the Na I D-lines. CCD échelle spectra (resolution 0.025 nm) obtained at Ritter Observatory between June 30 and Oct. 4 UT showed a single strong emission line at Hα, with essentially constant profile. On Oct. 8, the Hα line was markedly different, with an additional emission component present at −163 km s^{−1}. On the same night, the FWHM of the He I 587.5-nm emission line doubled, a P-Cyg absorption component of this line appeared at −200 km s^{−1}, and P-Cyg structure in the Na I D-lines developed. These sudden and unusual changes in spectral behavior have persisted and intensified through Oct. 14.”

V1357 CYGNI B. A. Harmon, G. J. Fishman, B. C. Rubin, and C. A. Wilson, Marshall

Space Flight Center, NASA; and W. S. Paciesas, University of Alabama at Huntsville, report for the Compton Observatory BATSE Team: “Hard x-ray emission from the black-hole candidate V1357 Cyg (Cygnus X-1), measured using the earth-occultation method, decreased gradually in the band 45-140 keV from a flux level of 0.12 to ∼ 0.05 photon cm^{−2} s^{−1} between early September and Oct. 10, and remained at a low flux level as of Oct. 17. The two flux levels correspond roughly to the γ_{2} (higher flux) and γ_{1} (lower flux) states observed in 1979-1980 by HEAO-3 (Ling et al. 1987, Ap.J. 321, L117). The September-October light curve is similar to that observed in May-June (IAUC 5813).”

R CORONAE BOREALIS Further visual magnitude estimates (cf. IAUC 5864), provided in part

by the AAVSO: Sept. 20.82 UT, 6.7 (G. Comello, Groningen, The Netherlands); 27.78, 6.7 (G. Zajacz, Debrecen, Hungary); Oct. 1.08, 6.6 (D. York, Abiquiu, NM); 4.08, 7.0 (York); 8.20, 6.6 (C. E. Spratt, Victoria, BC); 9.75, 7.5 (P. Schmeer, Bischmisheim, Germany); 11.75, 7.3 (L. Szentaskó, Budapest, Hungary); 16.81, 7.4 (A. Pereira, Cabo da Roca, Portugal); 19.80, 7.5 (Pereira).

1993 October 20 Daniel W. E. Green